# The InSAR absolute phase amid singularities

### v1

@@ 7.2.1 Single coherence observation
Two strategies for reconstructing the absolute phase from a single isolated coherence, in turn derived from one primary and one secondary observation, are commonly considered [6, 2]. In practice, they are commonly applied to multidimensional observations through averaging or selection of individual pixels.
@@
This strategy is viable whenever the absolute phase $\phi^{\mathrm{a}}$ is known to be defined for the given observation, thus ruling out an absolute phase singularity [6]. If it is known to be small in magnitude, $|\phi^{\mathrm{a}}|<\pi$, the absolute phase $\phi^{\mathrm{a}}=\varphi$, where the wrapped phase $\varphi$ is taken to be in the interval $(-\pi,\pi)$. Unwrapping along the (unobserved) absolute phase path $\Gamma^{\mathrm{a}}$ is unnecessary. More generally, the absolute phase may be reconstructed if it can be externally constrained to better than $2\pi$. Such a constraint will generally be easier to apply if path dependence can be ruled out.
@@
This strategy applies to point-like targets, for which absolute phase singularities cannot occur. The dependence of the absolute phase on the difference in range (or optical path) in (20) is proportional to the wavenumber $k$ and thus frequency, allowing for the estimation of the range difference within a wider ambiguity band than the first strategy [3]. This method, along with closely related radargrammetry approaches [30], is used in across-track interferometry, usually by averaging over multiple observations due to the high noise level [31]. However, the point-target assumption breaks down when the vertical extent approaches the height of ambiguity or, in differential InSAR, when the displacement variability is nonnegligible compared to the wavelength, or for dielectric changes [1, 32, 33]. Ionospheric phase contributions also invalidate (20). Having three or more frequencies may allow for estimating the ionospheric and the absolute, ionospherically corrected phase, or for detecting deviations from point-target behavior.
@@ 7.2.2 Samples along the absolute path
Serial observations along the absolute path facilitate absolute phase reconstruction. Assume that the ”‘absolute path dimension”’ is not included in the space $X_{\mathrm{d}}$. For instance, in the BorealScat data (Sec. 4.3.4), $X_{\mathrm{d}}$ encodes a spatial interferogram while the absolute path traverses time, so that integer phase offset estimation can be applied independently for each spatial pixel $x\in X_{\mathrm{d}}$. The discrete samples along the absolute path provide a sampled approximation to (12) is available. The denser the sample spacing, the easier the unwrapping and detection of zero or near-zero coherence magnitudes. In the BorealScat example, the 5-min spacing between successive secondary acquisitions facilitated reconstruction of (12) for the 1-day interferogram. However, determining unambiguously whether nearby points A and B in Fig. 5 differed in absolute phase by $\sim 2\pi$ would require infinitesimally dense samples. In across-track interferometry, multiple samples along the absolute path correspond to multiple baselines, the analysis of which resembles multifrequency processing for point-like targets [5].
For sparsely sampled observations along the absolute path, generalized closure phases can provide contextual information on absolute phase singularities [32, 23]. The simplest case is the closure phase [32]
@@
formed from three acquisitions at three positions along the absolute path, $s_{0}\leq s^{\prime}\leq s_{1}$. It depends on coherences formed from different primary acquisitions at $s_{0}$ and $s^{\prime}$ (the second argument of $\gamma$), in contrast to the absolute phase $\phi^{\mathrm{a}}(s_{1};s_{0})$ with primary $s_{0}$. This added complexity makes it difficult to constrain absolute phase reconstruction. However, large closure phase magnitudes indicate non-point-target behavior, occurring in scattering scenarios with dynamic vegetation or variable displacements [32, 33]. The insensitivity of closure phases to atmospheric phase screens makes them amenable to operational observational analyses.
@@ 8 Conclusion
This general absolute phase exhibits challenging properties. Its direct observation would require continuous measurements along the absolute path, while a regular interferogram only samples the endpoints. Fig. 1 illustrates the need for continuous observations: the absolute phase is path dependent, determined by intermediate states between $s=0$ and $s=1$ that discrete observations cannot capture. Sec. 6 shows that path dependence in configuration space $W^{+}$ is equivalent to the existence of unpatchable absolute phase singularities, with absolute phase jumps by nonzero integer multiples of $2\pi$. In observations, these absolute phase singularities can occur even for perfect coherence (Fig. 8), and they complicate observation-based reconstruction of the absolute phase.
The absolute phase usually needs to be reconstructed, using assumptions. In multidimensional interferograms, multifrequency methods (for point-like targets) or external information are commonly combined with phase unwrapping. Phase unwrapping can reconstruct the absolute phase up to a constant, provided certain conditions are met. Absolute phase singularities are a challenge as they generally leave no trace in the interferogram, and yet unwrapping paths are to avoid them. Sec. 7.1 re-interprets conditions in which the absolute phase can be reconstructed up to a constant through unwrapping [28], accounting for absolute phase singularities.
How useful a concept is the absolute phase, given these challenges? Path dependence implies that the absolute phase cannot be interpreted as a range difference, equivalent displacement or phase center position. Previous statements about the absolute phase’s essentiality for determining displacement or height from interferometric observations implicitly assumed point-like targets [5, 2], for which path dependence and absolute phase singularities do not occur. The general concept of the absolute phase is useful because it enables analysis of when point-target interpretations break down and what an unwrapped interferogram represents. In summary, the general definition and the theoretical analysis enable a comprehensive evaluation of InSAR products and processing chains.
Funding
Funded by NASA under awards 80NSSC25K7386, 80NSSC24K023.

### v3

@@ 7.2.1 Single coherence observation
Two strategies for reconstructing the absolute phase from a single isolated coherence, in turn derived from one primary and one secondary observation, are commonly considered [zebker98, rosen00]. In practice, they are commonly applied to multidimensional observations through averaging or selection of individual pixels.
@@
This strategy is viable whenever the absolute phase $\phi^{\mathrm{a}}$ is known to be defined for the given observation, thus ruling out an absolute phase singularity [zebker98]. If it is known to be small in magnitude, $|\phi^{\mathrm{a}}|<\pi$, the absolute phase $\phi^{\mathrm{a}}=\varphi$, where the wrapped phase $\varphi$ is taken to be in the interval $(-\pi,\pi)$. Unwrapping along the (unobserved) absolute phase path $\Gamma^{\mathrm{a}}$ is unnecessary. More generally, the absolute phase may be reconstructed if it can be externally constrained to better than $2\pi$. Such a constraint will generally be easier to apply if path dependence can be ruled out.
@@
This strategy applies to point-like targets that can be described by (20), for which absolute phase singularities cannot occur. The dependence of the absolute phase on the difference in range (or optical path) in (20) is proportional to the wavenumber $k$ and thus frequency, allowing for the estimation of the range difference within a wider ambiguity band than the first strategy [madsen93]. This method, along with radargrammetry, is used in across-track interferometry, usually by averaging over multiple observations due to the high noise level [brcic09]. However, (20) breaks down when the vertical extent approaches the height of ambiguity or, in differential InSAR, when the displacement variability is nonnegligible compared to the wavelength, or for dielectric changes [bamler98, dezan15, zwieback16]. Ionospheric phase contributions also invalidate (20). Three frequencies facilitate estimating the absolute ionospherically corrected phase and detecting deviations from point-target behavior.
@@ 7.2.2 Samples along the absolute path
Serial observations along the absolute path facilitate absolute phase reconstruction. Assume that the “absolute path dimension” is not included in the space $X_{\mathrm{d}}$. For instance, in the BorealScat data (Sec. 4.3.4), $X_{\mathrm{d}}$ encodes a spatial interferogram while the absolute path traverses time, so that integer phase offset estimation can be applied independently for each spatial pixel $x\in X_{\mathrm{d}}$. The discrete samples along the absolute path provide a sampled approximation to (12) is available. The denser the sample spacing, the easier the unwrapping and detection of zero or near-zero coherence magnitudes. In the BorealScat example, the 5-min spacing between successive secondary acquisitions facilitated reconstruction of (12) for the 1-day interferogram. However, determining unambiguously whether nearby points A and B in Fig. 5 differed in absolute phase by $\sim 2\pi$ would require infinitesimally dense samples. In across-track interferometry, multiple samples along the absolute path correspond to multiple baselines, the analysis of which resembles multifrequency processing for point-like targets [lombardini98].
For sparsely sampled observations along the absolute path, generalized closure phases can provide contextual information on absolute phase singularities [dezan15, zwieback24]. The simplest case is the closure phase [dezan15]
@@
formed from three acquisitions at three positions along the absolute path, $s_{0}\leq s^{\prime}\leq s_{1}$. It depends on coherences formed from different primary acquisitions at $s_{0}$ and $s^{\prime}$ (the second argument of $\gamma$), in contrast to the absolute phase $\phi^{\mathrm{a}}(s_{1};s_{0})$ with primary $s_{0}$. This added complexity makes it difficult to constrain absolute phase reconstruction. However, large closure phase magnitudes indicate a loss of correspondence between phase and range difference, as occurs for non-point-targets such as permafrost landforms with nonuniform displacement (Fig. 1) or vegetation [dezan15, zwieback16, ansari21]. The insensitivity of closure phases to atmospheric phase screens makes them amenable to operational observational analyses.
@@ 8 Conclusion
This general absolute phase exhibits challenging properties. Its direct observation would require continuous measurements along the absolute path, while a regular interferogram only samples the endpoints. Fig. 1 illustrates the need for continuous observations: the absolute phase is path dependent, determined by intermediate states between $s=0$ and $s=1$ that discrete observations cannot capture. Path dependence implies that the absolute phase is not proportional to a range difference. Sec. 6 shows that path dependence in the configuration space $W^{+}$ is equivalent to the existence of unpatchable absolute phase singularities, with absolute phase jumps by nonzero integer multiples of $2\pi$. In observations, these absolute phase singularities can occur even for perfect coherence (Fig. 9), and they complicate observation-based reconstruction of the absolute phase.
The absolute phase usually needs to be reconstructed using assumptions. For multidimensional interferograms, multifrequency methods (for point-like targets) or external information are commonly combined with phase unwrapping. Phase unwrapping can reconstruct the absolute phase up to a constant, provided certain conditions are met. Absolute phase singularities are a challenge as they generally leave no trace in the interferogram, and yet unwrapping paths are to avoid them. Sec. 7.1 re-interprets conditions in which the absolute phase can be reconstructed up to a constant through unwrapping [itoh82], accounting for absolute phase singularities. The reconstructed absolute phase may serve as input to displacement or elevation retrieval algorithms, though this assumes proportionality to range difference.
How useful is the absolute phase as defined here, given these challenges? Previously, the absolute phase formed the basis for displacement and elevation retrieval [madsen93, rosen00]. This relied on the assumption that it is proportional to range difference, an assumption valid for point targets. In general, however, this correspondence cannot hold. In differential interferometry the observational definition is the only way to assign a real quantity that is congruent with the wrapped phase and continuous (Sec. 4). The loss of correspondence to range difference severs the absolute phase’s direct connection to the equivalent displacement or phase center position practitioners seek to estimate. Downstream algorithms need to test whether the estimated absolute phase corresponds to a range difference, for instance through closure phases. Although the absolute phase need not have a simple physical interpretation, the observational definition and the theoretical tools introduced here enable principled evaluation of InSAR processing chains.